# RatBodyFormer: Rodent Body Surface from Keypoints

### v1

@@ 3.2 RatBodyFormer
The input parameters of RatBodyFormer are the keypoint coordinates $\mathbf{P}$, the individual-dependent scaling factors $\mathbf{C}=\{\mathbf{C}_{P},\mathbf{C}_{B}\}$, and translations $\mathbf{T}=\{\mathbf{T}_{P},\mathbf{T}_{B}\}$. We alternate between the optimization of RatBodyFomer and the individual-dependent parameters. The individual-dependent $\mathbf{C}$ and $\mathbf{T}$ are initialized by $\alpha$s and 00s, respectively. Here $\alpha$ is obtained as a result of aligning the canonical 3D body surface $\tilde{S}$ as mentioned before. At every epoch of the training, we optimize RatBodyFormer by minimizing $\mathcal{L}_{\mathrm{3D}}$, while keeping the individual-dependent parameters fixed. During this optimization, we refine $\mathbf{C}$ and $\mathbf{T}$ by optimizing $\mathcal{L}_{s}$ at every $T$ epochs, while keeping RatBodyFormer fixed. We use $T=50$ in our experiments.
@@
We estimate the scaling and translation parameters $\mathbf{C}$ and $\mathbf{T}$ of a new rat with inference-time optimization using the silhouette loss $\mathcal{L}_{s}$. The initial value $\alpha_{\mathrm{p}}$ for $\mathbf{C}_{P}$ and $\alpha_{\mathrm{B}}$ for $\mathbf{C}_{B}$ are manually set by the ratio of the length from the nose to the base of tail, and the ratio of the girth length, respectively. After optimizing $\mathbf{C}$ and $\mathbf{T}$, we regress the body surface point coordinates.
@@ 4 Experiments
We evaluate the effectiveness of RatBodyFormer with a number of experiments each focused on validating key properties of them. First, in Sec. 4.2, we evaluate the accuracy and generalizability of RatBodyFormer. Next, in Sec. 4.3, we demonstrate future 3D body surface prediction by forecasting with RatBodyFormer.
@@ 4.1 RatDome and RatDome Dataset
RatDome is equipped with 15 GoPro 10 cameras and 5 Intel L515 LiDAR-Camera devices. The LiDAR cameras are installed for potential ground truth measurements. We found that the depth captured with the LiDAR cameras were noisy and have too many holes for any direct measurement. We only use the depth images for D3 of Sec. 4.2 by combining it with shape-from-silhouette. GoPro captures 4K videos at 60Hz, and L515 captures 1080p videos at 30Hz and 1024$\times$768 depthmaps at 30Hz. We place a translucent acrylic tube of 300mm diameter and 5mm thick as a rat cage to avoid the rat from moving into the corners.
All the cameras are calibrated by capturing a chessboard moved around in the dome. The mean reprojection error after regular bundle adjustment was about 1.1 pixels [18]. Refraction by the acrylic tube is not modeled in the camera calibration. The cameras are temporally synchronized with flash light from a strobe.
@@ 4.2 RatBodyFormer
We evaluate the accuracy of RatBodyFormer in different scenarios. We first train and test the model using a single rat where the training and the testing sets share the same marker annotations (D1). We specifically evaluate the advantage of using semi-automatically annotated labels, the frames automatically annotated with RatBodyFormer itself (Sec. 3.2). Next, we combine two rats of different ages of weeks for training and testing (D2). This experiment quantifies the advantage of learning across individual rats. Finally, we use the entire RatDome Dataset, where we can use 7, 9, and 11 week-old annotations for training (D3), and use 5 and 14 week-old rats without markers to evaluate the generalizability of the proposed method.
@@
Fig. 6 shows the L2 error histograms for D1 between the predicted body surface 3D coordinates and the corresponding manually-annotated ground truths. This result demonstrates the advantage of using semi-automatically annotated labels as they improve the accuracy by about 1.7 mm. Fig. 5 shows qualitative results.
@@
Figure 6: L2 error histograms of D1. Each vertical bar indicates the mean L2 error of the histogram in the same color. Our semi-automatically annotation label improve average error by about 1.7 mm. “MA” and “SAA” mean “manually-annotated data”, and “semi-automatically annotated data”, respectively.
@@
Fig. 8 shows the results for D2. For this cross-individual experiments, the semi-automatically annotation labels also improve the accuracy. Fig. 7 shows qualitative results.
Figure 7: Qualitative results of D2. The left represents the results trained by only manually-annotated data (MA), and the right row shows the result of manually and semi-automatically annotated data (SAA). Our semi-automatically annotated labels also improve the quality in this scenario.
Figure 8: L2 error histograms of D2. Each vertical bar indicates the mean L2 error of the histogram in the same color. “MA” and “SAA” mean “manually-annotated data” and “semi-automatically annotated data”, respectively. Our semi-automatically annotated data improve L2 error by about 1.3mm.
@@
In this scenario, we train the model using 7, 9, and 11 week-old rats with the manually-annotated and semi-automatically annotated data and evaluate the accuracy on the 5 and 14 week-old rats without markers. For this, we define ground-truth surface points as the sum of the set of points captured by LiDAR and shape-from-silhouette. As many of the LiDAR points have missing values and errors, we add surface points obtained by shape-from-silhouette. We optimize individual-dependent parameters $\mathbf{C}$ and $\mathbf{T}$ with the silhouette loss Eq. 1. We obtain the mask images with SAM [24, 41]. The average estimation errors for 5 and 14 week-old rats are 3.98 mm and 6.39 mm, respectively. Note that we calculate only the dorsal side surface points because the ventral side sometimes occluded. Our model can estimate any shape of rat even if individuals not included in the training data. Fig. 9 shows the estimated body surface points. The results show that RatBodyFormer can be applied to a variety of rats ranging from 5 to 14 week-olds. This covers the age range typically used in biomedical and neuroscientific experiments.
@@
4.3 Pose Forecasting
In this section, we show that our surface model can predict future rat 3D body surface by just predicting detectable keypoint coordinates. As shown in Fig. 11, we train a transformer-based auto-regressive 3D keypoints prediction model. The input of this model is a sequence of keypoint 3D coordinates for past $\tau_{p}$ frames, and the output is the 3D keypoint coordinates of the next $\tau_{n}$ frames. We trained the model with 5, 7, 9, 11, and 14 week-old data for about 10000 frames in total and tested with 7 week-old data for about 500 frames.
Figure 10: Future rat body pose estimation. We take keypoint 3D coordinates for $\tau_{p}=3$ frames as input, and estimate $\tau_{n}=1,10,13$ future pose with the pose forecasting model and our RatBodyFormer. We can forecast future body poses by just predicting future keypoints. The accuracy degrades gracefully as we forecast further into the future.
Figure 11: Transformer encoder for keypoint forecasting. The input is a sequence of past concatenated $K=10$ keypoint 3D coordinates, and the output is 3D keypoint coordinates of the next frame. It predicts future keypoints’ coordinates auto-regressively.
Figure 12: Histogram of 3D keypoint prediction errors. Each vertical bar indicates the mean L2 error of the histogram in the same color. As the peak average error at $\tau_{p}=1$ is about 3mm, we can predict 3D keypoint coordinates in most cases. Since errors accumulate, the larger $\tau_{n}$ is, the bigger its error is.
Fig. 12 shows the histogram of L2 errors in the predicted keypoints at $\tau_{p}=3$ and $\tau_{n}=1,3,5$ frames. At $\tau_{n}=1$, our model can predict the future keypoint positions in most cases accurately, except for when the rat goes under rapid movements, such as suddenly changing directions. We leave invesitigation of more sophisticated methods for complex motion prediction as future work. Fig. 10 shows predictions of the body surface points by RatBodyFormer. Given predicted 3D keypoint coordinates in the future frames, RatBodyFormer can infer the 3D body surface coordinates from them. This is possible as RatBodyFormer effectively reduces the necessary points for forecasting to those that are well-defined. We believe this forecasting scheme can benefit a wide-range of scientific experiments, especially that concern multiple rats.
@@ 5 Conclusion
This paper proposed the RatDome Dataset and RatBodyFormer. The RatDome Dataset is the first-of-its-kind dataset which provides 3D body surface coordinates with temporally-consistent annotations from 7, 9, and 11 week-old rats. RatBodyFormer models the highly nonrigid deformation of the rat body and can regress the body surface 3D coordinates from its keypoint 3D coordinates. Experimental results demonstrated that RatBodyFomer can generalize to model the body surface points of 5 and 14 week-old rats which are not included in the training dataset.
We believe that our RatDome Dataset and RatBodyFormer collectively serve a novel, sound foundation for autonomous rat behavior analysis and will likely have far-reaching implications for biomedical and neuroscientific research.
@@ Appendix A Implementation Details of Canonical Body Surface
Appendix B Evaluations with Other Data Splits
To demonstrate the generalizability of our RatBodyFormer, we further evaluate the accuracy of RatBodyFormer with other data splits. Tab. A shows the dataset splits of D1 and D2 scenarios. The splits “D1” and “D2” in Table 1 of the main paper appear as D1-a and D2-a in this table, respectively.
@@ Appendix B Evaluations with Other Data Splits
Figure B: L2 error histograms of D1-b, D1-c, and D1-d. Each vertical bar indicates the mean L2 error of the histogram in the same color. “MA” and “SAA” denote “manually-annotated data”, and “semi-automatically annotated data”, respectively. Our semi-automatically annotated data improves accuracy by 1.6 mm to 2.3 mm.
@@
Fig. D shows the error histogram of D2-b and Fig. E shows qualitative results. These results demonstrate quantitatively and qualitatively that our semi-automatically annotated data improves accuracy regardless of the dataset splits.
Figure D: L2 error histogram of D2-b. Each vertical bar indicates the mean L2 error of the histogram in the same color. ‘MA” and “SAA” represent “manually-annotated data”, and “semi-automatically annotated data”, respectively. The semi-automatically annotated data improves average error by about 1.6 mm. ‘
Figure E: Qualitative results of D2-b. We show the results of RatBodyFormer trained with only manually-annotated data (MA) in the left, and that trained with both manually-annotated and semi-automatically annotated data (SAA) in the right. Semi-automatically annotated data clearly improves the body surface estimation.
Appendix C Ablation Study
C.1 Individual-dependent Parameters
This section evaluates the effect of the individual-dependent parameters, i.e., point-wise scaling parameters $\mathbf{C}$ and translation parameters $\mathbf{T}$ for the training time and inference time. Fig. F evaluates the contribution of optimizing the individual-dependent parameters at training time using D2-a and D2-b splits. For the case of training without optimizing individual-dependent parameters, we set $\mathbf{C}=1$ and $\mathbf{T}=0$. These results show that optimizing the individual-dependent parameters at training time improves the estimation accuracy.
Fig. G shows the effect of optimizing the individual-dependent parameters at inference time using the D3 split in Table 1 of the main text. In this split, RatBodyFormer is trained using 7-, 9-, and 11-week-old rats with manually-annotated and semi-automatically annotated data, and evaluated with 5- and 14-week-old rats using 3D points obtained by LiDAR and shape-from-silhouette as described in the main text. For the results of inference without optimizing individual-dependent parameters, we set $\mathbf{C}=1$ and $\mathbf{T}=0$. We can observe that the inference time optimization clearly improves the estimation accuracy.

### v3

@@ 3.2 RatBodyFormer
The input parameters of RatBodyFormer are the keypoint coordinates $\mathbf{P}$, the individual-dependent scaling factors $\mathbf{C}=\{\mathbf{C}_{P},\mathbf{C}_{B}\}$, and translations $\mathbf{T}=\{\mathbf{T}_{P},\mathbf{T}_{B}\}$. We alternate between the optimization of RatBodyFormer and the individual-dependent parameters. The individual-dependent $\mathbf{C}$ and $\mathbf{T}$ are initialized by $\alpha$s and 00s, respectively. Here $\alpha$ is obtained as a result of aligning the canonical 3D body surface $\tilde{S}$ as mentioned before. At every epoch of the training, we optimize RatBodyFormer by minimizing $\mathcal{L}_{\mathrm{3D}}$ and $\mathcal{L}_{\mathrm{p}}$, while keeping the individual-dependent parameters fixed. During this optimization, we refine $\mathbf{C}$ and $\mathbf{T}$ by optimizing $\mathcal{L}_{\mathrm{s}}$ and $\mathcal{L}_{\mathrm{p}}$ at every $T$ epochs, while keeping RatBodyFormer fixed. We use $T=50$ in our experiments.
@@
We estimate the scaling and translation parameters $\mathbf{C}$ and $\mathbf{T}$ of a new rat with inference-time optimization using $\mathcal{L}_{\mathrm{s}}$ and $\mathcal{L}_{\mathrm{p}}$. The initial value $\alpha_{\mathrm{p}}$ for $\mathbf{C}_{P}$ and $\alpha_{\mathrm{B}}$ for $\mathbf{C}_{B}$ are manually set by the ratio of the length from the nose to the base of the tail, and the ratio of the girth length, respectively. After optimizing $\mathbf{C}$ and $\mathbf{T}$, we regress the body surface point coordinates.
@@ 4 Experiments
We evaluate the effectiveness of RatBodyFormer with a number of experiments each focused on validating key properties of it. In Sec. 4.2, we evaluate the accuracy and generalization capability of RatBodyFormer. We also compare with MAMMAL [2].
@@ 4.1 RatDome and RatDome Dataset
RatDome is equipped with 15 GoPro 10 cameras and 5 Intel L515 LiDAR-Camera devices. The LiDAR cameras are installed for potential ground truth measurements of markerless rats. We found that the depth captured with the LiDAR cameras was noisy and had holes for any direct measurement. We only use the depth images for D3 of Sec. 4.2 by combining it with shape-from-silhouette. GoPro captures 4K videos at 60Hz, and L515 captures 1080p videos at 30Hz and 1024$\times$768 depthmaps at 30Hz. We place a translucent acrylic tube of 300 mm diameter and 5 mm thick as a rat cage to avoid the rat from moving into the corners.
All the cameras are calibrated by capturing a chessboard moved around in the dome. The mean reprojection error after regular bundle adjustment was about 1.1 pixels [22]. Refraction by the acrylic tube is not modeled in the camera calibration. The cameras are temporally synchronized with flash light from a strobe. The flash light can potentially be observed over two frames leading to a 1/60s misalignment.
@@ 4.2 RatBodyFormer
We evaluate the accuracy of RatBodyFormer in different scenarios. We first train and test the model using a single rat where the training and the testing sets share the same marker annotations (D1). We specifically evaluate the advantage of using semi-automatically annotated labels, the frames automatically annotated with RatBodyFormer itself (Sec. 3.2). Next, we combine two rats of different ages of weeks for training and testing (D2). This experiment also quantifies the advantage of using semi-automatically annotated labels in the case of combining two rats. Finally, we use the entire RatDome Dataset, where we can use 7, 9, and 11 week-old annotations for training (D3), and use 5, 7, 9, 11, and 14 week-old rats without markers to evaluate the generalization capability of the proposed method.
@@
Fig. 5 shows the L2 error histograms for D1 between the predicted body surface 3D coordinates and the corresponding manually-annotated ground truths. This result demonstrates the advantage of using semi-automatically annotated labels as they improve the accuracy by about 0.9 mm. Fig. 4 shows qualitative results.
@@
Figure 5: L2 error histograms of D1. Each vertical bar indicates the mean L2 error of the histogram in the same color. Our semi-automatically annotation label improve average error by about 0.9 mm. “MA” and “SAA” mean “manually-annotated data”, and “semi-automatically annotated data”, respectively.
@@
The average estimation error of D2 is 7.3 mm when trained using manually annotated labels, and 6.6 mm when trained with both manually and semi-automatically annotated labels. For this case of combining two rats of different ages, the semi-automatically annotation labels also improve the accuracy. This result also shows that RatBodyFormer can estimate the body surface with an error comparable to the diameter of a bead (5.5 mm) for rats of different ages.
@@
In this scenario, we train the model using 7, 9, and 11 week-old rats with the manually-annotated and semi-automatically annotated data and evaluate the accuracy on the 5, 7, 9, 11, 14 week-old rats without markers. For this, we define ground-truth surface points as the sum of the set of points captured by LiDAR and shape-from-silhouette. As the LiDAR points have missing values and errors, we add surface points obtained by shape-from-silhouette. We optimize individual-dependent parameters $\mathbf{C}$ and $\mathbf{T}$ with the silhouette loss and the position consistency loss for 5 and 14 week-old rats. We obtain the mask images with SAM [33, 50]. The average estimation errors for 5 and 14 week-old rats are 1.80 mm and 2.40 mm, respectively. On the other hand, the average estimation errors for 7, 9, and 11 week-old rats are 2.69 mm, 1.97 mm, and 2.43 mm, respectively. Note that we calculate only the dorsal surface points because the ventral side is sometimes occluded. Even for individuals with body shapes not included in the training data, RatBodyFormer can estimate the body surface with an error comparable to that of individuals with body shapes present in the training data. Fig. 6 shows the estimated body surface points. The results show that RatBodyFormer can be applied to a variety of rats ranging from 5 to 14 week-olds. This covers the age range typically used in biomedical and neuroscientific experiments.
@@
Comparison with MAMMAL
We compare the accuracy of RatbodyFormer with that of MAMMAL [2] for the body surface marker points using D2 in Tab. 1. MAMMAL uses the mouse mesh model scanned in [8]. As the body size of a mouse is different from a rat, we scale the virtual mouse mesh model [8] to match the scale of the beaded rat with a similarity transform using 8 keypoints, i.e., nose, right and left ears, front paws, back paws, and the base of the tail, in their rest poses. Then, we ARAP deform the surface bead points of the rat to the mouse mesh, using the 8 keypoints as its hard constraints and surface points as its soft constraints. The target position used in the soft constraint for each surface point is chosen as the mesh vertex of the mouse model closest from the surface point at each ARAP iteration. After the alignment, we select the mesh vertices of the mouse model that are closest to the surface markers of the beaded rat as the positions of the beads in the mouse model.
Fig. 7 shows the L2 error histograms of MAMMAL and RatBodyFormer. RatBodyFormer achieves lower L2 errors. Note that we calculate errors only for the dorsal surface points because the ventral side is occluded in the rest pose. Fig. 8 shows qualitative results. Our RatBodyFormer achieves higher accuracy by a large margin.
Figure 7: L2 error histogram for RatBodyFormer and MAMMAL. Each vertical bar shows the mean L2 error of each histogram in the same color. Our method achieves higher accuracy.
Figure 8: Qualitative comparison between RatBodyFormer (right) and MAMMAL (left). RatBodyFormer achieves smaller displacements (black arrows) of the estimated marker positions from the ground truth. The projections of the 3D mesh by MAMMAL visualize its inaccurate 3D shape fitting.
5 Application
The dense body surface recovered by RatBodyFormer can enable a variety of downstream applications towards rich rat behavior analysis. Here, we demonstrate one such application of deriving an animatable 3D rat model. We refer to this model as GaussianRat as we anchor 3D Gaussian splatting to the estimated body surface. We define the 3D body surface mesh by using the surface points estimated by RatBodyFormer for the torso and the points from shape-from-silhouette for the face. RatBodyFormer lets us drive the 3D body surface mesh through the placement of 3D keypoints.
Inspired by human head avatars [51, 49, 54], we attach 3D Gaussians to each mesh face and optimize their parameters. Each Gaussian has six properties, location $\bm{u}$, scale $s$, rotation $\bm{r}$, opacity $\sigma$, color $\bm{h}$, and attached mesh face ID $k$. We define each Gaussian position $\bm{u}$ by the normalized barycentric coordinates $x,y,z$ and the pose dependent distance $l$ ($l=l^{\prime}+w\mathbf{P}$) in the normal direction of the face. Here, $l^{\prime}$, $w$, $\mathbf{P}$ denotes the pose-independent distance, weights, and keypoint 3D coordinates, respectively. We optimized the parameters except for the face ID $k$, by minimizing the L1 loss and the D-SSIM loss between the original and the reconstructed images [51, 49, 54].
Fig. 9 shows the rendered GaussianRat and Fig. 10 shows the animated GaussianRat. Please refer to the appendix for more details. GaussianRat not only enables photo-realistic body surface estimation, but also is animatable through 3D keypoint locations. We believe GaussianRat can serve as a useful tool for analysis-by-synthesis to decipher complex rat behaviors, including those in complex interactions with other individuals and can also directly be used as virtual stimuli.
Figure 9: GaussianRat enables photo-realistic rendering of the rat body surface (bottom left) from detected keypoints (orange, left) and recovered body surface (top right).
Figure 10: Frames of an animated GaussianRat.
@@ 6 Conclusion
This paper proposed the RatDome Dataset and RatBodyFormer. The RatDome Dataset is the first-of-its-kind dataset which provides 3D body surface coordinates with temporally-consistent annotations from 7, 9, and 11 week-old rats. RatBodyFormer models the highly nonrigid deformation of the rat body and can regress the body surface 3D coordinates from its keypoint 3D coordinates. Experimental results demonstrate its accuracy, generalization capability, and the potential of wide application.
Since each keypoint set corresponds uniquely to a body surface point set in the RatDome Dataset, RatBodyFormer takes only keypoints as input. But in situations not included in the RatDome Dataset, such as complex interactions, a single keypoint set may not uniquely determine a body surface point set. We plan to address this by incorporating explicit visual cues of the shape of the body surface, such as silhouettes, in our future work. We believe that our RatDome Dataset and RatBodyFormer collectively serve a novel, sound foundation for autonomous rat behavior analysis and will likely have far-reaching implications for neuroscientific research.
@@ Appendix A Implementation Details of Canonical Body Surface
Appendix B Implementation Details of Position Consistency Loss
RatBodyFormer estimates the body surface points of all annotated body surface points. But a single input keypoint set from a rat has a ground-truth annotation only for the body surface points of the same rat. A position consistency loss $\mathcal{L}_{\mathrm{p}}$ is introduced to improve the position consistency of all the estimated body surface points. For each body surface point $j$, we first define three neighboring points $j_{1},j_{2},j_{3}$ in the reference pose. Then, we calculate the barycentric coordinates $(\alpha_{j_{1}},\alpha_{j_{2}},\alpha_{j_{3}})$ of the foot of the perpendicular dropped from $\tilde{b}_{j}$ to the triangle formed by $\tilde{b}_{j_{1}},\tilde{b}_{j_{2}},\tilde{b}_{j_{3}}$. Here, $\tilde{B}=\{~{}\tilde{b}_{j}\}_{1}^{N}$ denotes the body surface points in the reference pose. The position consistency loss $\mathcal{L}_{\mathrm{p}}$ is calculated as:
$\mathcal{L}_{\mathrm{p}}=\sum_{j}\left\{\begin{smallmatrix}0&\text{if $(\alpha_{j_{1}}^{\prime},\alpha_{j_{2}}^{\prime},\alpha_{j_{3}}^{\prime}>0,\alpha_{j_{1}}^{\prime}+\alpha_{j_{2}}^{\prime}+\alpha_{j_{3}}^{\prime}=1)$},\\ \Delta_{j}&\text{else},\end{smallmatrix}\right.$ (1)
where $\Delta_{j}=\|(\alpha_{j_{1}}^{\prime}-\alpha_{j_{1}})b_{j_{1}}^{\prime}+(\alpha_{j_{2}}^{\prime}-\alpha_{j_{2}})b_{j_{2}}^{\prime}+(\alpha_{j_{3}}^{\prime}-\alpha_{j_{3}})b_{j_{3}}^{\prime}\|_{2}$, $\mathbf{B}^{\prime}=\{b^{\prime}_{j}\}_{1}^{N}$ is the estimated body surface points, and $(a_{j_{1}}^{\prime},a_{j_{2}}^{\prime},a_{j_{3}}^{\prime})$ are the barycentric coordinates of the foot of the perpendicular dropped from ${b}^{\prime}_{j}$ to the triangle formed by ${b}^{\prime}_{j_{1}},{b}^{\prime}_{j_{2}},{b}^{\prime}_{j_{3}}$. We experimentally show the effectiveness of a position consistency loss in E.1.
Appendix C Evaluations with Other Data Splits
To demonstrate the generalization capability of our RatBodyFormer, we further evaluate the accuracy of RatBodyFormer with other data splits. Tab. A shows the dataset splits of D1 and D2 scenarios. The splits “D1” and “D2” in Table 1 of the main paper appear as D1-a and D2-a in this table, respectively.
@@ Appendix C Evaluations with Other Data Splits
Figure B: L2 error histograms of D1-b, D1-c, and D1-d. Each vertical bar indicates the mean L2 error of the histogram in the same color. “MA” and “SAA” denote “manually-annotated data”, and “semi-automatically annotated data”, respectively. Our semi-automatically annotated data improves accuracy by 0.7 mm to 2.2 mm.
@@
Fig. D shows the error histogram of D2-b. This results demonstrate that our semi-automatically annotated data improves accuracy regardless of the dataset splits.
Figure D: L2 error histogram of D2-b. Each vertical bar indicates the mean L2 error of the histogram in the same color. ‘MA” and “SAA” represent “manually-annotated data”, and “semi-automatically annotated data”, respectively. The semi-automatically annotated data improves average error by about 1.7 mm.
Appendix D Evaluations of GausianRat
In this section, we show the effectiveness of GaussianRat. To evaluate the photometric consistency of GaussianRat, we compare SSIM, PSNR, and LPIPS of GaussianRat to those of one scene optimization. We have synchronized 15 multiview cameras, and use 9 for testing and 6 for training. Tab. B shows the quantitative comparison. In Fig. E, we show the qualitative results of a novel view of GaussianRat and one scene optimization. We can reconstruct the appearance consistently compared to optimizing one scene individually, when rendered from a novel view.
Method SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$
GaussianRat $\bm{0.559}$ $\bm{16.348}$ $\bm{0.229}$
One scene optimization 0.233 13.210 0.331
Table B: Quantitative comparison. GaussianRat outperforms one scene optimization on all the metrics.
Figure E: Qualitative results of novel-view synthesis. Compared with optimizing one scene individually, GaussianRat demonstrates photometric consistency.
To evaluate the geometric-consistency, we compare the transition of each Gaussian center across 15 frames with Dynamic 3D Gaussians [40], which tracks the Gaussians over time. As shown in F, GaussianRat demonstrates higher geometric consistency.
Figure F: The Gaussian centers of Dynamic 3D Gaussians (top) and GaussianRat (bottom). The same Gaussians are represented with the same color across all frames. GaussianRat maintains consistency in the positions of the Gaussians on the body surface over time, whereas Dynamic 3D Gaussians lacks such consistency.
Appendix E Ablation Study
E.1 Position Consistency Loss
This section evaluates the effect of a position consistency loss $\mathcal{L}_{\mathrm{p}}$ using D2-a and D2-b in Tab. A. In this dataset, the body surface points which RatBodyFormer estimates are defined as the union of annotated body surface points on 7- and 11-week old rats. Fig. G qualitatively shows the effect of a position consistency loss. Fig. H shows L2 error histograms of estimated body surface points with and without a position consistency loss. Note that we can only calculate the error of own body surface points, and we cannot calculate it of the other individual body surface points. A position consistency loss improves the consistency of estimated body surface point positions without substantially degrading the estimation accuracy of own body surface points.
Figure G: Qualitative results of the effect of a position consistency loss. We manually define the mesh using the body surface points in the reference pose. These body surface points are the union of the body surface points of 7- and 11-week old rats. In the case without a position consistency loss, some estimated body surface points lacks position consistency (red circle).